OpenStudy (anonymous):

Why does sqrt(L^2-((1/2)L)^2) = (sqrt(3)/2) x L ?

OpenStudy (anonymous):

\[\sqrt{L^{2} - (\frac{ 1}{ 2 }}L)^{2} = \frac{ \sqrt{3} }{ 2 }L\]\] Can someone show me the steps they would take to go from the left to the right please :)

OpenStudy (anonymous):

 just if your interested, this is where im getting it from, solving for H

OpenStudy (turingtest):

\[\sqrt{L^2-(\frac12L)^2}=\sqrt{L^2-\frac14L^2}=\sqrt{\frac44L^2-\frac14L^2}\]\[=\sqrt{\frac34L^2}=\frac{\sqrt3}2L\]

OpenStudy (anonymous):

thanks, ahhh L^2 = 1L^2 = (4/4)L^2 yeahhh thats the bit. Thanks alot